Circular No. 5924

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993N IN UGC 5695 A. V. Filippenko and T. Matheson, University of California at Berkeley,

write: “CCD spectrograms (range 310-1000 nm) obtained on Jan. 7 and 19 UT with the 3-m Shane reflector at Lick Observatory show that the type-II supernova 1993N is indeed similar to SN 1988Z, as previously suspected from the early-time spectrum (cf. IAUC 5788). The dominant feature is Hα emission, with FWHM ∼ 1800 km/s and its centroid blueshifted by 500-600 km/s relative to the parent galaxy’s systemic velocity of 2900 km/s. Weak Hβ and Hγ emission are also visible above the faint continuum. By analogy with SN 1988Z and its close cousin SN 1986J, the ejecta of SN 1993N are interacting with a dense circumstellar wind. It is probable that SN 1993N will become a luminous x-ray and radio source within the next few years. Having less than half the distance of SN 1988Z, and having been discovered far earlier in its evolution than SN 1986J, SN 1993N offers an unprecedented opportunity to learn more about this rare but interesting subclass of type-II supernovae.”

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) D. Jewitt and J. Chen, University of Hawaii, report images of this

comet obtained on Jan. 14 UT using the UH 2.2-m telescope on Mauna Kea. Under 0′′.8 seeing, the ‘string of pearls’ morphology of the comet persists, with significant changes in the relative brightnesses of some components. The most notable change, however, is the near-disappearance of the dust sheet previously observed in extension to the north-east and south-west of the line of nuclei. In addition, individual tails from the separate nuclei have become more apparent, indicating persistent release of dust from these nuclei. The longest tail (from nucleus 7) is in excess of 1′ along p.a. 295^{o}.

COMET KUSHIDA (1994a) Total visual magnitude estimates: Jan. 12.25 UT, 10.8 (A. Hale, near

Orogrande, NM, 0.20-m reflector); 12.93, 12.3 (K. Sárneczky, Veresegyház, Hungary, 0.33-m reflector); 13.96, 11.6 (T. Vanmunster, Landen, Belgium, 0.35-m reflector); 16.25, 11.1 (G. Kronk, Troy, IL, 0.33-m reflector); 18.03, 10.6 (G. Comello, Groningen, The Netherlands, 0.20-m reflector).

1994 January 20 Daniel W. E. Green